# Mosaic IT: Enhancing Instruction Tuning with Data Mosaics

### v1

@@ 4.1 Implementation Details
The experiments are conducted on three base pre-trained models: Llama2-7B, Llama2-13B [46], and Mistral-7B [19]. The training datasets and detailed training configurations are introduced in Appendix B.
@@ 4.2 Evaluation Metrics
Pair-wise Comparison by using powerful LLMs like GPT-4 is recently widely accepted and becoming a common practice [46, 7, 11, 35, 6]. The evaluation of responses from LLMs, especially in open-domain contexts where definitive ground truth is hard to establish, continues to be an intricate and evolving research domain. Recent studies, however, have indicated a notable alignment between GPT-4’s performance evaluations and human assessments [66, 29, 42], thereby establishing a credible foundation for this evaluative methodology. We adopted test instruction sets from WizardLM [54], comprising 218 diverse, human-curated instructions for pair-wise comparison. We directly follow the evaluation framework proposed by [3, 27], which evaluates responses on a scale spanning from 1 to 10 across multiple dimensions. To further address positional bias, as discussed by [22, 48], the comparison is conducted in two distinct sequences, LLM1’s response first and then LLM2’s response first, ensuring a fair assessment of model performance. Evaluation outcomes are categorized into ’win-tie-loss’ for each instruction. The detailed evaluation prompt is provided in Appendix D.
Human Evaluation is further implemented to substantiate the superiority of our approach based on the WizardLM test set. The test set contains $100$ samples randomly sampled from the original WizardLM test set. Three human evaluators were tasked with comparing the outputs generated by the models under consideration, using the same criteria as in the previous pairwise evaluation. Each evaluator was presented with three response options: Win, Tie, and Loss. The final outcomes were determined by a majority vote. More evaluation metrics including Open LLM Leaderboard, Alpaca-Eval Leaderboard, MT-Bench and IF Eval are introduced in Appendix B.
@@ 5 Experimental Results
5.1 Main Results
In this section, we present the evaluation results comparing our methods with the baseline methods on Three baseline models (Mistral-7B [19], Llama2-7B [46], Llama2-13B) and Three instruction tuning datasets (Alpaca-GPT4 [40], Alpaca [44], WizardLM-70k [54]), on Two general evaluation settings (Pair-Wise Comparison and Open LLM leaderboard) described above, as shown in the Table 1. Pair-Wise Winning Score indicates the result directly comparing our models with the corresponding baseline models, which is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$. These values that are greater than $1.0$ represent better responses generated by our models. The performances on the Huggingface Open LLM Leaderboard are also presented, and we bold the greater average values for each comparison. The consistent outperforming results on different base models and datasets represent the effectiveness and robustness of our methods. Moreover, comparing the results on different base models, the performance growths of our method on Mistral-7B and Llama2-13B models are greater than the Llama2-7B models, indicating that our method works better on relatively better base models. We hypothesize it is because our method largely increases the complexity of instructions by proposing a higher level of meta-instructions while keeping the quality of original responses unchanged, making it hard for weaker models to learn.
To better understand how our method improves the instruction-following abilities of LLMs, we further compare the performance on other Three benchmarks for fine-grained analysis based on the Mistral-7B base model with two datasets as shown in Table 2. On the MT-Bench, the 1-round scores of our method are higher while the 2-round scores are slightly lower, indicating that our method mainly improves the response quality for single-round conversations, which is reasonable as the meta instructions only focus on single-round formats. On the Alpaca Eval benchmark, our method has a consistent improvement with or without the Length Control (LC), indicating that the improvement of response qualities does not directly originate from the length of responses. On the IF Eval benchmark, our method consistently improves the performances on all 4 different settings, both Prompt-level and Instruction-level, both Strict version and Loose version. Compared with the previous benchmarks, IF Eval mainly focuses on the format-following ability of LLMs. The consistent improvement in this benchmark represents that our method not only improves the response qualities of the LLMs but also improves their controllability regarding formats. Given that our method is a cost-free augmentation technique that does not rely on any additional models, the observed improvements are remarkable.
@@ 5.1 Main Results
Dataset Method MT-Bench $\uparrow$ Alpaca Eval 2 $\uparrow$ IF Eval $\uparrow$
1-round 2-round Rate Rate (LC) Prompt (S) Inst (S) Prompt (L) Inst (L)
Alpaca-GPT4 Baseline 6.44 5.26 3.98 7.28 32.53 42.93 35.86 45.92
@@
Table 2: The performance comparison on the MT-Bench, Alpaca Eval, and IF Eval Benchmarks. Rate(LC) in Alpaca Eval represents the length-controlled win rates. In IF Eval, Prompt and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions.
@@ 5.2 Ablation studies
In this section, extensive experiments are conducted on Mistral-7B using with the Alpaca-GPT4 dataset to verify the effectiveness of our method. We utilize Pair-wise comparison for evaluation.
@@
(a) Ablation on Mosaic-IT strategies.
@@
(b) Ablation on the Max Number of Instructions.
Ablation on Mosaic Strategies is presented in Table 3(a). Primary represents the Primary Mosaic Strategy. The winning score of this setting is greater than $1.0$, indicating a better performance compared with the baseline method. This comparison directly verifies the effectiveness of the idea of introducing multiple instructions during training, which complicates the instructions at no cost and improves the instruction-following ability of LLMs. Format represents the Format Strategy. Although the winning score is only slightly greater than the naive version, this version makes it possible for LLMs to follow the customized user-defined formats, indicating great potential for the controllability of LLMs. Moreover, the format version can be easily used with other types of meta instructions, showing great extensibility. Permute represents the Permute Strategy that builds on the Format Strategy with a probability of $\frac{1}{2}$, similar to Maskout. Permute/Maskout represents our default setting, where the Permute or Maskout Strategies are utilized together with the Format Strategie with a probability of $\frac{1}{3}$. All these $3$ settings show higher performance than the format version, indicating the effectiveness of Auxiliary Mosaic Strategies which define more complicated meta instructions.
@@
Table 4: Ablation on the Distribution of Number of Instructions, including the pair-wise comparison values (left) and the visualization of distribution comparisons (right). The detailed numbers of instructions for different settings can be shown in Appendix A. “Mix $\leq 5$” represents the percentage of samples with the number of instructions less or equal to $5$.
Ablation on the Distribution of Number of Instructions is presented in Table 4, including the pair-wise comparison values (left) and detailed number distribution comparisons (right), which aims at identifying how this count distribution affects the performance of our method. The detailed distribution formula, and data counts are provided in the Appendix A. Fix represents the setting where all the overall instructions are concatenated with a fixed number of instructions, which we set as $10$ unless the overall instructions exceed the max length limit. Exponential represents the setting where the number of instructions is sampled following the exponential distribution. Under these two settings, less than $3\%$ of the overall instructions are concatenated by less or equal to $5$ original instructions. The lack of few-instruction concatenated samples negatively affects the LLMs’ ability to follow the single instruction, which is employed by most of the existing evaluation methods, leading to worse performances. Pareto, Log-normal, and Logistic represents the corresponding distribution that are utilized for sampling. Different from the above two settings, approximately $10\%$ of the overall instructions are composed of fewer original instructions, thus ensuring the LLMs are trained with samples with sufficiently diverse lengths, resulting in optimal performances. Uniform is our default setting, representing using the uniform distribution where different numbers are sampled evenly. In this situation, the LLMs are trained with samples with the most diverse lengths, thus avoiding the LLMs overfit to simple lengthy responses.
@@ 6 Further Discussion
Figure 2: The training loss curve comparisons between the original instruction tuning process and our Mosaic-IT with w datasets on (a) Mistral-7B, (b) Llama2-7B, and (c) Llama2-13B. Two conclusions can be drawn from the comparisons: (1) Much fewer training iterations are required in our method, indicating the more efficient training process; (2) The “stair-like” loss curves for the original training process indicate a potential memorizing effects, while our loss curves are more smooth. All the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc.
6.1 Memorizing
@@
In the original instruction tuning process, each data sample will be trained several times for LLMs without many changes to the instructions and responses. Though effective, this training process poses risks to the potential memorizing effects on training samples, which can be partially indicated by the “stair-like” training loss curves as shown in Figure 2. In the figure, all the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc. For the baseline methods, the training loss hardly decreases within each epoch of training but drops dramatically when the LLMs meet the same training samples again, which indicates a potential memorizing effect of training samples and potential overfitting. However, when utilizing our method, the random mosaics of original instructions with diverse and complex meta-instructions largely diversify the overall training instructions. Although each original data sample will still be seen by LLMs several times during training, the overall context varies dramatically as each original sample is only an atomic element of the overall mosaic sample, indicating that there will be no identical overall instructions during the whole training process. Thus this augmentation largely alleviates the potential memorizing and overfitting problems as shown in the figure, where the training loss decreases smoothly, representing the gradual learning process.
6.2 Efficiency
@@
Table 5: The training time comparison of different settings, and the pair-wise winning scores are also provided for better illustration. “Uni-2” represents uniform distribution with max count as $2$. Mosaic-IT can reduce the training time to approximately $16\%$ to $25\%$ while achieving better performance.
One of the benefits of our method is the efficiency of the training process. Given an existing dataset, our mosaic processes largely decrease the number of total overall instructions and the total number of gradient descents, leading to a reduction in the training process. The detailed comparison is shown in Table 5, which is based on the Mixtral-7B model on the Alpaca-GPT4 dataset. The time is calculated based on four NVIDIA A100 Graphic Cards. As shown, our method greatly decreases the training time to approximately $16\%$ to $25\%$ while achieving better performances, especially when there are mosaic samples with larger permutation counts.
6.3 Limitation
The potential limitations of our work: (1) Currently, three Auxiliary Mosaic Strategies with corresponding high-level rules are proposed and utilized in our method, however, we believe more strategies and predefined rules can be further introduced. (2) The optimal distribution of the number of instructions for the mosaic process still needs further justification in future studies. (3) It is unknown whether the inclusion of extra models or careful curation/selection of instructions for concatenation will further improve the performance of Mosaic-IT largely.

### v2

@@ 3.1 Implementation Details
The experiments are conducted on: Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024). The training datasets include Alpaca Taori et al. (2023), Alpaca-GPT4 Peng et al. (2023), WizardLM Xu et al. (2023), Vicuna 1M Zheng et al. (2024a), and Magpie Xu et al. (2024b) datasets. Due to the extremely large size of Vicuna 1M and Magpie, $300$k instances are randomly sampled for our experiments. The detailed description of datasets and the training configurations are introduced in Appendix B.
@@ 3.2 Evaluation Metrics
We utilize five automatic evaluation metrics, including (i) LLM-based Pair-wise Comparison, (ii) Open LLM leaderboard, (iii) MT-Bench, (iv) Alpaca Eval, and (v) IF Eval, and (vi) Human evaluation to verify the effectiveness of our method. They are widely accepted evaluation metrics for measuring LLMs’ instruction-following capabilities. The introductions of the five automatic evaluation metrics are provided in Appendix B.
Human Evaluation is further implemented to substantiate the superiority of our approach based on the WizardLM test set. The test set contains $100$ samples randomly sampled from the original WizardLM test set. Three human evaluators were tasked with comparing the outputs generated by the models under consideration, using the same criteria as in the previous pairwise evaluation. Each evaluator was presented with three response options: Win, Tie, and Loss. The final outcomes were determined by a majority vote.
@@ 4 Experimental Results
Table 2: The performance comparison on the MT-Bench, Alpaca Eval, and IF Eval Benchmarks. Rate(LC) in Alpaca Eval represents the length-controlled win rates. In IF Eval, Prompt, and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions.
Model Dataset Method MT-Bench $\uparrow$ Alpaca Eval 2 $\uparrow$ IF Eval $\uparrow$
1-round 2-round Rate (LC) Rate Prompt (S) Inst (S) Prompt (L) Inst (L)
Mistral 7B Alpaca-GPT4 Baseline 6.44 5.26 3.98 7.28 32.53 42.93 35.86 45.92
@@
Table 3: The performance comparison on more model families and datasets on all five automatic evaluation metrics. In IF Eval, P and I represent Prompt-level and Instruction-level accuracy.
Model Dataset Method Pair-wise $\uparrow$ Open LLM $\uparrow$ Alpaca Eval 2    $\uparrow$ MT-Bench $\uparrow$ IF Eval $\uparrow$
Score Average Rate (LC) Rate 1-round 2-round P(L) I(L)
Llama-3-8B Vicuna Baseline 1.000 52.51 2.15 1.36 6.70 5.06 21.26 33.45
Mosaic-IT 1.234 55.62 3.09 2.05 6.85 5.40 31.42 45.56
Magpie Baseline 1.000 56.15 9.22 13.74 8.10 7.08 35.67 47.72
Mosaic-IT 1.133 60.13 12.23 16.05 8.36 7.49 40.67 52.76
Phi-3 Vicuna Baseline 1.000 62.06 4.20 2.74 5.34 4.18 30.50 43.17
Mosaic-IT 1.083 62.30 5.95 3.83 5.89 4.53 32.35 41.85
Magpie Baseline 1.000 62.90 13.82 17.68 7.78 6.42 44.36 55.52
Mosaic-IT 1.014 63.54 14.04 17.67 7.89 6.16 50.83 62.35
Gemma2-2B Vicuna Baseline 1.000 48.90 1.72 1.31 6.69 5.25 23.66 35.61
Mosaic-IT 1.266 51.31 1.90 1.38 6.93 5.26 24.03 36.93
Magpie Baseline 1.000 46.37 5.35 7.77 4.57 3.23 21.81 32.49
Mosaic-IT 1.032 48.36 5.66 8.54 5.16 3.96 22.18 34.77
4.1 Main Results
In this section, we present the evaluation results comparing our methods with the baseline methods on several baseline models (Mistral-7B Jiang et al. (2023), Llama2-7B Touvron et al. (2023b), Llama2-13B) and instruction tuning datasets (Alpaca-GPT4 Peng et al. (2023), Alpaca Taori et al. (2023), WizardLM-70k Xu et al. (2023)), on Two general evaluation settings (Pair-Wise Comparison and Open LLM leaderboard) described above, as shown in the Table 1. Pair-wise Winning Score indicates the result directly comparing our models with the corresponding baseline models, which is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$. These values that are greater than $1.0$ represent better responses generated by our models. The performances on the Huggingface Open LLM Leaderboard are also presented, and we bold the greater average values for each comparison. The consistent outperforming results on different base models and datasets represent the effectiveness and robustness of our methods.
To better understand how our method improves the instruction-following abilities of LLMs, we further compare the performance on other Three benchmarks for fine-grained analysis based on the Mistral-7B base model with two datasets as shown in Table 2. On the MT-Bench, the 1-round scores of our method are higher, indicating that our method mainly improves the response quality for single-round conversations, which is reasonable as the meta instructions only focus on single-round formats. On the Alpaca Eval benchmark, our method has a consistent improvement with or without the Length Control (LC), indicating that the improvement of response qualities does not directly originate from the length of responses. On the IF Eval benchmark, our method consistently improves the performances on all 4 different settings, both Prompt-level and Instruction-level, both Strict version and Loose version. Compared with the previous benchmarks, IF Eval mainly focuses on the constraint-following ability of LLMs. The consistent improvement in this benchmark represents that our method not only improves the response qualities of the LLMs but also improves their controllability regarding formats. Given that our method is a cost-free augmentation technique that does not rely on any additional models, the observed improvements are remarkable.
Moreover, to further verify the effectiveness of our method, more experiments on different model families and data families are conducted, as shown in Table 3, including Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024) models on Vicuna 1M Zheng et al. (2024a), and Magpie Xu et al. (2024b) datasets. For these two datasets, $300$k data are randomly sampled to verify the scalability of our method when dealing with large amounts of instruction-tuning data. The performances of our models consistently outperform the baseline models across different model families and data sources, ranging from diverse data qualities.
@@ 4.2 Ablation studies
In this section, extensive ablation experiments are conducted on Mistral-7B using with the Alpaca-GPT4 dataset to verify our method. We utilize Pair-wise comparison for evaluation.
(a) Ablation on Mosaic-IT strategies.
@@
(b) Ablation on the Max Number of Instructions.
@@
Ablation on Mosaic Strategies is presented in Table 4(a). Primary represents the Primary Mosaic Strategy. The winning score of this setting is greater than $1.0$, indicating a better performance compared with the baseline method. This comparison directly verifies the effectiveness of the idea of introducing multiple instructions during training, which complicates the instructions at no cost and improves the instruction-following ability of LLMs. Format represents the Format Strategy. Although the winning score is only slightly greater than the naive version, this version makes it possible for LLMs to follow the customized user-defined formats, indicating great potential for the controllability of LLMs. Moreover, the format version can be easily used with other types of meta instructions, showing great extensibility. Permute represents the Permute Strategy that builds on the Format Strategy with a probability of $\nicefrac{{1}}{{2}}$, similar to Maskout. Permute/Maskout represents our default setting, where the Permute or Maskout Strategies are utilized together with the Format Strategie with a probability of $\nicefrac{{1}}{{3}}$. All these $3$ settings show higher performance than the format version, indicating the effectiveness of Advanced Mosaic Strategies which define more complicated meta instructions.
Table 5: Ablation on the Distribution of Number of Instructions. The distribution formula and data counts for different settings are shown in Appendix A. “Mix $\leq 5$” represents the percentage of samples with the number of instructions less or equal to $5$.
@@
Figure 3: Ablation on the Distribution of Number of Instructions, the visualization of distribution comparisons.
Ablation on the Distribution of Number of Instructions is presented, including the pair-wise comparison values in Table 5 and detailed number distribution comparisons in Figure 3, which aims at identifying how this count distribution affects the performance of our method. The detailed distribution formula and data counts are provided in the Appendix A. Fix represents the setting where all the overall instructions are concatenated with a fixed number of instructions, which we set as $10$ unless the overall instructions exceed the max length limit. Exponential represents the setting where the number of instructions is sampled following the exponential distribution. Under these two settings, less than $3\%$ of the overall instructions are concatenated by less or equal to $5$ original instructions. The lack of few-instruction concatenated samples negatively affects the LLMs’ ability to follow the single instruction, which is employed by most of the existing evaluation methods, leading to worse performances. Pareto, Log-normal, and Logistic represents the corresponding distribution that are utilized for sampling. Different from the above two settings, approximately $10\%$ of the overall instructions are composed of fewer original instructions, thus ensuring the LLMs are trained with samples with sufficiently diverse lengths, resulting in optimal performances. Uniform is our default setting, representing using the uniform distribution where different numbers are sampled evenly. In this situation, the LLMs are trained with samples with the most diverse lengths, thus avoiding the LLMs overfit to simple lengthy responses.
@@ 5 Further Discussion
5.1 Preliminary Experiments: Performance Degradation
The motivation of our Mosaic-IT is also rooted in the observation that when handling multiple instructions simultaneously, a performance degradation will incurred for even strong LLMs like GPT-4-turbo. While LLMs generally perform well when responding to single instructions, their capability to follow multiple instructions at once tends to decline noticeably. BatchPrompt has shown the uncertainty when LLMs are requested to answer multiple formatted questions at one time. Moreover, in some cases, e.g for general open-domain instructions, LLMs might directly ignore some of the instructions, especially when the LLMs are required to respond to the instructions in a random pre-defined order, which is exactly simulating our Permute strategy.
@@ 5.1 Preliminary Experiments: Performance Degradation
Figure 4: The training loss curve comparisons between the original instruction tuning process and our Mosaic-IT with w datasets on (a) Mistral-7B, (b) Llama2-7B, and (c) Llama2-13B. The “stair-like” loss curves for the original training process indicate potential memorizing effects, while our loss curves are smoother. All the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc.
To quantitatively analyze this phenomenon, experiments using GPT-3.5-turbo and GPT-4-turbo are conducted on the WizardLM test set. Specifically, we compare the models’ performance when responding to multiple instructions concurrently versus responding to a single instruction at each time, by utilizing LLM-based Pair-Wise comparison, as shown in Table 6. All the win rates are lower than $1.0$, demonstrating a clear and significant reduction in response quality when these models are required to respond to multiple instructions at one time. Moreover, the possibility of missing instructions (Miss Rate) increases further when they are required to respond to the instructions in a predefined random order rather than a sequential order. These results clearly demonstrate the difficulties of following several instructions at a time and why it can be regarded as a higher level of instruction-following capability.
Table 6: Pair-wise win rate of performances when responding to multiple instructions concurrently versus responding to a single instruction each time, and miss rate when responding to multiple instructions concurrently. “3 Instructions” represents the setting where 3 random instructions are concatenated together for inference. “Sequential” and “Random” represents the setting where the models are asked to respond to each instruction sequentially, or in a random pre-defined order.
Pair-Wise (Multi vs. Single) 3 Instructions 5 Instructions 7 Instructions
Win Rate $\uparrow$ Miss Rate $\downarrow$ Win Rate $\uparrow$ Miss Rate $\downarrow$ Win Rate $\uparrow$ Miss Rate $\downarrow$
GPT-3.5-turbo (Sequential) 0.357 0.014 0.336 0.055 0.303 0.064
GPT-3.5-turbo (Random) 0.315 0.124 0.330 0.156 0.198 0.312
GPT-4-turbo (Sequential) 0.176 0.000 0.137 0.000 0.140 0.000
GPT-4-turbo (Random) 0.139 0.000 0.153 0.014 0.101 0.005
5.2 Mosaic: Alleviating Memorizing
@@
In the original instruction tuning process, each data sample will be trained several times for LLMs without changes to the instructions and responses. This training process poses risks to the potential memorizing effects on training samples, which can be partially indicated by the “stair-like” training loss curves as shown in Figure 4. In the figure, all the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc. For the baseline methods, the training loss hardly decreases within each epoch of training but drops dramatically when the LLMs meet the same training samples again, which indicates a potential memorizing effect of training samples and potential overfitting. However, when utilizing our method, the random mosaics of original instructions with diverse and complex meta-instructions largely diversify the overall training instructions. Although each original data sample will still be seen by LLMs several times during training, the overall context varies dramatically as each original sample is only an atomic element of the overall mosaic sample, indicating that there will be no identical overall instructions during the whole training process. Thus this augmentation largely alleviates the potential memorizing and overfitting problems as shown in the figure, where the training loss decreases smoothly, representing the gradual learning process.
5.3 Mosaic: Improving Efficiency
Table 7: The training time comparison of different settings, and the pair-wise winning scores are also provided for better illustration. “Uni-2” represents uniform distribution with max count as $2$. Mosaic-IT reduces the training time to $16\%-25\%$ while achieving better performance.
@@ 5.3 Mosaic: Improving Efficiency
One of the benefits of our method is the efficiency of the training process. Given an existing dataset, our mosaic processes largely decrease the number of total overall instructions and the total number of gradient descents, leading to a reduction in the training process. The detailed comparison is shown in Table 7, which is based on the Mistral-7B model on the Alpaca-GPT4 dataset. The time is calculated based on four NVIDIA A100 Graphic Cards. As shown, our method greatly decreases the training time to approximately $16\%$ to $25\%$ while achieving better performances, especially when there are mosaic samples with larger permutation counts.
5.4 Mosaic: Why It Works?
The effect of our method is aligned with the Dense and Aligned Captions Doveh et al. (2023) used in VL Models, which utilizes denser captions to promote the VL models. Different from all previous methods which require LLMs to generate the whole response conditioned on the whole instruction, our method forces LLMs to generate responses selectively conditioned on some critical parts of the overall dense instruction. Especially in advanced strategies, LLMs are required to generate responses conditioned not only on sequential parts of the instruction, but diverse and randomized segments of it, which is defined by meta-instruction.
Compared to original settings, our setting requires LLMs to first understand the formats and orders defined in the meta-instructions, then adapt to conditioning on different parts of the instructions when generating responses. This process forces LLMs to develop a more comprehensive understanding of context, prioritize various pieces of information, and manage complex dependencies between instructions, thus improving instruction-following performance. Moreover, the entire process is data- and model-agnostic, ensuring the generalizability of our method.
5.5 Limitation and Future Directions
The potential limitations of our work: (1) Currently, three Advanced Mosaic Strategies with corresponding high-level rules are proposed and utilized in our method, however, we believe more strategies and predefined rules can be further introduced. (2) The optimal distribution of the number of instructions for the mosaic process still needs further justification in future studies. (3) It is unknown whether the inclusion of extra models or careful curation/selection of instructions for concatenation will further improve the performance of Mosaic-IT largely.